if cos x= -1/4 and pi/2 < x < pi, find the value of sin(x/2)
Answer by Tutorteddy.Com(12)

cos x = 1 - 2sin^2(x/2)
=> sin^2(x/2) = (1 - cosx)/2 = (1 + 1/4)/2 = 5/8
=> sinx/2 = Sqrt(5/8) = [sqrt(5/2)]2
The negative value is discarded as pi/2 < x < pi and sinx is positive in this quadrant.